OpenStudy (anonymous):

Write a situation that will use the equation 21=2x+7

OpenStudy (texaschic101):

hmmm...oh I hate these. 2x + 7 = 21 2x = 21 - 7 2x = 14 x = 7 you go into a store and you see some things you want, but you only have $21. you want (something that costs $7-- just make up something) and you also want some candy. The candy (x) costs $2. How many pieces of candy can you buy without going over your budget of $21 ? By solving the equation, you find out that you can buy 7 pieces. Does that sound to stupid ??

OpenStudy (anonymous):

No lol it's okay

OpenStudy (texaschic101):

It doesn't even have to be candy...just something that costs $2 a piece.

OpenStudy (texaschic101):

you need any more help ?

OpenStudy (anonymous):

Yeah but i'm going to open up a different question give me a second